Q5. (5 pts). Suppose the current spot rate is $ 1.55/£ on the first of January. By year's end, the US CPI is expected to climb from 144 to 150 and the UK CPI is expected to climb from 120 to 130. According to PPP, what is the expected spot rate on December 31? (Show your calculations!)
Answer #1

As per Purchasing Power Parity, the price of a commodity in two countries is the same

Spot Rate = $1.55/Pound

% Inflation in US = (150-144)/144 = 4.1667%

% inflation rate in UK = (130-120)/120 = 8.3333%

Expected Spot Rate on Dec 31 = Spot Rate(1+Inflation Rate Dollar)/(1+Inflation Rate Pound)

= 1.55(1+4.1667%)/(1+8.3333%)

= $1.4904/Pound